### v1

@@ 3.3.2 Self-Distillation
For Medusa-1, this dataset is sufficient for training Medusa heads. However, for Medusa-2, we observe that solely using this dataset for training the backbone and Medusa heads usually leads to a lower generation quality. In fact, even without training Medusa heads, training the backbone model with this dataset will lead to performance degradation. This suggests that we also need to use the original model’s probability prediction instead of using the ground truth token as the label for the backbone model, similar to classic knowledge distillation works (Kim and Rush, 2016). Concretely, the loss for the backbone model is:
@@
Figure 3: Visualization of a sparse tree setting for Medusa-2 Vicuna-7B. The tree has depth 4 which indicates 4 Medusa heads involved in calculation. Each node indicates a token from a top-k prediction of a Medusa head, and the edges show the connections between them. The red lines highlight the path that correctly predicts the future tokens.
@@ 3.3.3 Searching for the Optimized Tree Construction
In Section 3.1.2, we present the simplest way to construct the tree structure by taking the Cartesian product. However, with a fixed number of total nodes in the tree, a regular tree structure may not be the best choice. Intuitively, those candidates composed of the top predictions of different heads may have different accuracies. Therefore, we can leverage an estimation of the accuracy to construct the tree structure.
Specifically, we can use a calibration dataset and calculate the accuracies of the top predictions of different heads. Let $a_{k}^{(i)}$ denote the accuracy of the $i$-th top prediction of the $k$-th head. Assuming the accuracies are independent, we can estimate the accuracy of a candidate sequence composed by the top $\left[i_{1},i_{2},\cdots,i_{k}\right]$ predictions of different heads as $\prod_{j=1}^{k}a_{j}^{(i_{j})}$. Let $I$ denote the set of all possible combinations of $\left[i_{1},i_{2},\cdots,i_{k}\right]$ and each element of $I$ can be mapped to a node of the tree (not only leaf nodes but all nodes are included). Then, the expectation of the acceptance length of a candidate sequence is:
@@
Thinking about building a tree by adding nodes one by one, the contribution of a new node to the expectation is exactly the accuracy associated with the node. Therefore, we can greedily add nodes to the tree by choosing the node that is connected to the current tree and has the highest accuracy. This process can be repeated until the total number of nodes reaches the desired number. In this way, we can construct a tree structure that maximizes the expectation of the acceptance length.
Fig. 3 illustrates the structure of a sparsely constructed tree for the Medusa-2 Vicuna-7B model. This tree structure extends four levels deep, indicating the engagement of four Medusa heads in the computation. The tree is initially formed through a Cartesian product approach and subsequently refined by pruning based on the statistical expectations of the top-k predictions from each Medusa head measured on the Alpaca-eval dataset Dubois et al. (2023). The tree’s lean towards the left visually represents the algorithm’s preference for nodes with higher probabilities on each head.
@@ 4 Experiments
In this section, we present two sets of experiments to demonstrate the effectiveness of Medusa under different settings. First, we evaluate Medusa on the Vicuna-7B and 13B models (Chiang et al., 2023) to show the performance of Medusa-1 and Medusa-2. Second, we evaluate Medusa on the Vicuna-33B and Zephyr-7B (Tunstall et al., 2023) models to study the effectiveness of self-distillation because for Vicuna-33B model, the training dataset is not publicly available, and for Zephyr-7B model, the model is trained with RLHF.
4.0.1 Shared Settings
@@
4.1.1 Experimental Setup
We use the Vicuna model class (Chiang et al., 2023), which encompasses chat models of varying sizes (7B, 13B, 33B) that are fine-tuned from the Llama model (Touvron et al., 2023). Among them, the 7B and 13B models are trained on the ShareGPT (ShareGPT, 2023) dataset, while the 33B model is an experimental model and is trained on a private dataset. In this section, we use the ShareGPT dataset to train the Medusa heads on the 7B and 13B models for $2$ epochs. We use the v1.5 version of Vicuna models, which are fine-tuned from Llama-2 models with sequence length 4096. We use a global batch size of $64$ and a peak learning rate of $5e^{-4}$ for the backbone and $2e^{-3}$ for Medusa heads and warmup for $40$ steps. We use $4$-bit quantized backbone models for both models. We first train the models with Medusa-1 and use these trained models as initialization to train Medusa-2. We employ QLoRA for Medusa-2 and the $\lambda_{0}$ in Eq. (2) is set to be $0.2$.
Figure 4: Left: Speed comparison of baseline, Medusa-1 and Medusa-2 on Vicuna-7B/13B. Medusa-1 achieves more than 2$\times$ wall-time speedup compared to the baseline implementation while Medusa-2 further improves the speedup by a significant margin. Right: Detailed speedup performance of Vicuna-7B on 8 categories from MT-Bench.
4.1.2 Results
We collect the results and show them in Fig. 4. The baseline is the vanilla Huggingface implementation. In Fig. 3(a), we can see that for the 7B models, Medusa-1 and Medusa-2 configurations lead to a significant increase in speed, measuring in tokens processed per second. Medusa-1 shows a 2.18$\times$ speedup, while Medusa-2 further improves this to a 2.83$\times$. When applied to the larger 13B model, Medusa-1 results in a 2.33$\times$ speed increase, while Medusa-2 maintains a similar performance gain of 2.83$\times$ over the baseline. We also plot the speedup per category for Medusa-2 Vicuna-7B model. We observe that the “Coding” category benefits from a 3.29$\times$ speedup, suggesting that Medusa is particularly effective for tasks in this domain. This points to a significant potential for optimizing coding LLMs, widely used in software development and other programming-related tasks. The “Extraction” category shows the highest speedup at 3.62$\times$, indicating that this task is highly optimized by the Medusa. Overall, the results suggest that the Medusa significantly enhances inference speed across different model sizes and tasks.
@@ 4.2 Case Study: Training with Self-Distillation on Vicuna-33B and Zephyr-7B
4.2.1 Experimental Setup
In this case study, we focus on the cases where self-distillation is needed. We use the Vicuna-33B model (Chiang et al., 2023) and the Zephyr-7B model (Tunstall et al., 2023) as examples. Following the procedure described in Section 3.3.2, we first generate the datasets with some seed prompts. We use ShareGPT (ShareGPT, 2023) and UltraChat (Ding et al., 2023) as the seed datasets and collect a dataset at about $100k$ samples for both cases. Interestingly, we find that the Zephyr model can continue to generate multiple rounds of conversation with a single prompt, which makes it easy to collect a large dataset. For Vicuna-33B, we generate the multi-turn conversations by iteratively feeding the prompts from each multi-turn seed conversation. Both models are trained with sequence length $2048$ and batch size $128$. We use Medusa-2 for both models and instead of using a two-stage training procedure, we use a sine schedule for the $\theta_{0}$ to gradually increase the value to its peak at the end of the training, we find this approach is equally effective. We set the peak learning rate of the backbone LoRA adapter to be $1e^{-4}$ and the warmup steps to be $20$. Since the self-distillation loss is relatively small, we set the $\lambda_{0}$ in Eq. (2) to be $0.01$.
4.2.2 Results
Table 1 complements these findings by comparing various Medusa-2 models in terms of their acceleration rate, overhead, and quality on MT-Bench. Notably, while the Medusa-2 Vicuna-33B model shows a lower acceleration rate, it maintains a comparable quality. We hypothesize that this is due to a mismatch between the hidden training dataset and the dataset we used for self-distillation.
Figure 5: Speedup of various models with Medusa-2. Medusa-2 shows significant speed improvement over all the models, while models trained with self-distillation have weaker speedup due to the trade-off between preserving model quality and boosting model speed.
@@ 4.2.2 Results
Table 1: Comparison of various Medusa-2 models. The quality denotes the average scores on the MT-Bench benchmark (Zheng et al., 2023). Medusa-2 achieves promising acceleration rate with mild overhead and preserves the model quality.
Figure 6: Evaluation of additional length introduced by trees. Left: The acceleration rate for randomly sampled dense tree settings (blue dots) and optimized sparse tree settings (red stars). Right: The speed (tokens/s) for both settings. The trend lines indicate that while the acceptance rate remains relatively stable for sparse trees, there is a notable decrease in speed as the additional length increases.
@@ 4.3.1 Configuration of Tree Attention
The ablation study of tree attention is conducted on the writing and roleplay categories from the MT-Bench dataset (Zheng et al., 2023) using Medusa-2 Vicuna-7B. We target to depict tree attention’s motivation and its performance.
Fig. 5(a) compares the acceleration rate of randomly sampled dense tree configurations (Section. 3.1.2, depicted by blue dots) against optimized sparse tree settings (Section. 3.3.3, shown with red stars). The sparse tree configuration with 64 nodes shows a better acceleration rate than the dense tree settings with 256 nodes. Fig. 5(b) presents the speed for both dense and sparse tree settings. The trend observed here indicates a notable decrease in speed as the additional length increases. This suggests that while sparse tree attention is beneficial for maintaining a steady acceleration rate, it comes with the trade-off of reduced speed, particularly at higher additional lengths.
The observed decline in speed is attributed to the increased overhead introduced by the hardware architecture. While a more complex tree can improve the acceleration rate, it does so at the cost of speed due to the hardware-imposed overhead.
Figure 7: Performance comparison of Medusa using proposed typical sampling. The plot illustrates the acceleration rate (Acc. Rate) and averaged scores on the “Writing” and “Roleplay” (MT benchmark) with a fixed temperature of 0.7 for 3 different settings: greedy sampling with Medusa, random sampling, and typical sampling under different thresholds. The model is fully fine-tuned Vicuna-7B.
@@
Table 2: Comparison of Different Settings Vicuna-7B. Quality is obtained by evaluating models on MT-Bench.
4.3.2 Thresholds of Typical Acceptance
The thresholds of typical acceptance are studied on the “Writing” and “Roleplay” categories from the MT-Bench dataset (Zheng et al., 2023) using Medusa-2 Vicuna 7B. Utilizing the Vicuna 7B model, we aligned our methodology with the approach delineated by  Hewitt et al. (2022) setting the $\alpha=\sqrt{\epsilon}$. Fig. 7 presents a comparative analysis of our model’s performance across various sampling settings. These settings range from a threshold $\epsilon$ starting at 0.01 and incrementally increasing to 0.25 in steps of 0.01. Our observations indicate a discernible trade-off: as $\epsilon$ increases, there is an elevation in quality at the expense of a reduced acceleration rate. Furthermore, for tasks demanding creativity, it is noted that the default random sampling surpasses greedy sampling in performance, and the proposed typical sampling is comparable with random sampling when $\epsilon$ increases.
@@ 4.3.3 Effectiveness of Two-stage Fine-tuning
We examine the performance differences between two fine-tuning strategies for the Vicuna-7B model in Table 2. We provided the comparison of directly fine-tuning the model with the Medusa heads vs. Medusa-2 that involves two-stage fine-tuning described in Section 3.2.2. The findings indicate that implementing our Medusa-2 for fine-tuning maintains the model’s quality and concurrently improves the speedup versus Medusa-1.
4.4 Discussion
In conclusion, we have proposed Medusa, a novel method to accelerate large language model inference by equipping models with additional predictive decoding heads. Medusa allows models to generate multiple tokens per step, overcoming the bottleneck of sequential auto-regressive decoding. We have demonstrated two procedures, Medusa-1 and Medusa-2, for efficiently training these extra heads while preserving model performance. Experiments on models of varying sizes and training methods show consistent speedups of 2.3-3.6$\times$ on single prompt inference across different prompt types and models.
Key advantages of Medusa include its simplicity, parameter efficiency, and ease of integration into existing systems. By building on top of speculative decoding concepts, Medusa avoids the need for specialized draft models. The typical acceptance scheme also removes complications from rejection sampling while still providing reasonable outputs. Finally, the fine-tuning procedures ensure high-quality generations without affecting the performance of the original model.

### v2

@@ 2.3.2 Self-Distillation
For Medusa-1, this dataset is sufficient for training Medusa heads. However, for Medusa-2, we observe that solely using this dataset for training the backbone and Medusa heads usually leads to a lower generation quality. In fact, even without training Medusa heads, training the backbone model with this dataset will lead to performance degradation. This suggests that we also need to use the original model’s probability prediction instead of using the ground truth token as the label for the backbone model, similar to classic knowledge distillation works (Kim & Rush, 2016). Concretely, the loss for the backbone model is:
@@ 2.3.3 Searching for the Optimized Tree Construction
In Section 2.1.2, we present the simplest way to construct the tree structure by taking the Cartesian product. However, with a fixed budget for the number of total nodes in the tree, a regular tree structure may not be the best choice. Intuitively, those candidates composed of the top predictions of different heads may have different accuracies. Therefore, we can leverage an estimation of the accuracy to construct the tree structure.
Specifically, we can use a calibration dataset and calculate the accuracies of the top predictions of different heads. Let $a_{k}^{(i)}$ denote the accuracy of the $i$-th top prediction of the $k$-th head [^2] Here, the accuracy is defined for the single top $i$-th token, i.e., this accuracy is equal to top-$i$ accuracy minus top-$(i-1)$ accuracy.. Assuming the accuracies are independent, we can estimate the accuracy of a candidate sequence composed by the top $\left[i_{1},i_{2},\cdots,i_{k}\right]$ predictions of different heads as $\prod_{j=1}^{k}a_{j}^{(i_{j})}$. Let $I$ denote the set of all possible combinations of $\left[i_{1},i_{2},\cdots,i_{k}\right]$ and each element of $I$ can be mapped to a node of the tree (not only leaf nodes but all nodes are included). Then, the expectation of the acceptance length of a candidate sequence is:
@@
Thinking about building a tree by adding nodes one by one, the contribution of a new node to the expectation is exactly the accuracy associated with the node. Therefore, we can greedily add nodes to the tree by choosing the node that is connected to the current tree and has the highest accuracy. This process can be repeated until the total number of nodes reaches the desired number. In this way, we can construct a tree that maximizes the expectation of the acceptance length. Further details can be found in Appendix C.
Figure 3: Left: Speed comparison of baseline, Medusa-1 and Medusa-2 on Vicuna-7B/13B. Medusa-1 achieves more than 2$\times$ wall-time speedup compared to the baseline implementation while Medusa-2 further improves the speedup by a significant margin. Right: Detailed speedup performance of Vicuna-7B with Medusa-2 on 8 categories from MT-Bench.
@@ 3 Experiments
In this section, we present experiments to demonstrate the effectiveness of Medusa under different settings. First, we evaluate Medusa on the Vicuna-7B and 13B models (Chiang et al., 2023) to show the performance of Medusa-1 and Medusa-2. Then, we assess our method using the Vicuna-33B and Zephyr-7B models to demonstrate self-distillation’s viability in scenarios where direct access to the fine-tuning recipe is unavailable, as with Vicuna-33B, and in models like Zephyr-7B that employ Reinforcement Learning from Human Feedback (RLHF). The evaluation is conducted on MT-Bench (Zheng et al., 2023), a multi-turn, conversational-format benchmark. Detailed settings can be found in Appendix B.
@@
Experimental Setup. We use the Vicuna model class (Chiang et al., 2023), which encompasses chat models of varying sizes (7B, 13B, 33B) that are fine-tuned from the Llama model (Touvron et al., 2023). Among them, the 7B and 13B models are trained on the ShareGPT (ShareGPT, 2023) dataset, while the 33B model is an experimental model and is trained on a private dataset. In this section, we use the ShareGPT dataset to train the Medusa heads on the 7B and 13B models for $2$ epochs. We use the v1.5 version of Vicuna models, which are fine-tuned from Llama-2 models with sequence length 4096.
Results. We collect the results and show them in Fig. 3. The baseline is the default Huggingface implementation. In Fig. 3(a), we can see that for the 7B models, Medusa-1 and Medusa-2 configurations lead to a significant increase in speed, measuring in tokens processed per second. Medusa-1 shows a 2.18$\times$ speedup, while Medusa-2 further improves this to a 2.83$\times$. When applied to the larger 13B model, Medusa-1 results in a 2.33$\times$ speed increase, while Medusa-2 maintains a similar performance gain of 2.83$\times$ over the baseline. We also plot the speedup per category for Medusa-2 Vicuna-7B model. We observe that the coding category benefits from a 3.29$\times$ speedup, suggesting that Medusa is particularly effective for tasks in this domain. This points to a significant potential for optimizing coding LLMs, which are widely used in software development and other programming-related tasks. The “Extraction” category shows the highest speedup at 3.62$\times$, indicating that this task is highly optimized by the Medusa. Overall, the results suggest that the Medusa significantly enhances inference speed across different model sizes and tasks.
@@ 3.2 Case Study: Training with Self-Distillation on Vicuna-33B and Zephyr-7B
Experimental Setup. In this case study, we focus on the cases where self-distillation is needed. We use the Vicuna-33B model (Chiang et al., 2023) and the Zephyr-7B model (Tunstall et al., 2023) as examples. Following the procedure described in Section 2.3.2, we first generate the datasets with some seed prompts. We use ShareGPT (ShareGPT, 2023) and UltraChat (Ding et al., 2023) as the seed datasets and collect a dataset at about $100k$ samples for both cases. Interestingly, we find that the Zephyr model can continue to generate multiple rounds of conversation with a single prompt, which makes it easy to collect a large dataset. For Vicuna-33B, we generate the multi-turn conversations by iteratively feeding the prompts from each multi-turn seed conversation using random sampling with temperature 0.3. Both models are trained with sequence length $2048$ and batch size $128$.
@@
$S_{\textnormal{SpecDecoding}}$ 1.47 - 1.56 1.60
$S_{\textsc{Medusa}}$ 2.83 2.66 2.83 2.35
Table 1: Comparison of various Medusa-2 models. The first section reports the details of Medusa-2, including accelerate rate, overhead, and quality that denoted the average scores on the MT-Bench compared to the original models. The second section lists the speedup ($S$) of SpecDecoding and Medusa, respectively.
Figure 4: Effectiveness of numbers of candidate tokens for decoding introduced by trees (default number of candidate token for decoding is 1 when using KV cache). Left: The acceleration rate for randomly sampled dense tree settings (blue dots) and optimized sparse tree settings (red stars). Right: The speed (tokens/s) for both settings. The trend lines indicate that while the acceleration rate remains relatively stable for sparse trees, there is a notable decrease in speed as the candidate tokens increases.
Results. Table 1 complements these findings by comparing various Medusa-2 models in terms of their acceleration rate, overhead, and quality on MT-Bench with GPT-4 acting as the evaluator to assign performance scores ranging from 0 to 10. We report the quality differences of Medusa compared to the original model. Notably, while the Medusa-2 Vicuna-33B model shows a lower acceleration rate, it maintains a comparable quality. We hypothesize that this is due to a mismatch between the hidden training dataset and the dataset we used for self-distillation. Hence, the model’s generation quality can be well aligned by self-distillation while Medusa heads learn distribution from the self-distillation that potentially shifts from the training set. In our study, we also applied speculative decoding (Chen et al., 2023; Leviathan et al., 2022) to the Vicuna lineup using open-source draft models (details can be found in Appendix D).
@@ 3.3.1 Configuration of Tree Attention
The study of tree attention is conducted on the writing and roleplay categories from the MT-Bench dataset using Medusa-2 Vicuna-7B. We target to depict tree attention’s motivation and its performance.
Fig. 4(a) compares the acceleration rate of randomly sampled dense tree configurations (Section. 2.1.2, depicted by blue dots) against optimized sparse tree settings (Section. 2.3.3, shown with red stars). The sparse tree configuration with 64 nodes shows a better acceleration rate than the dense tree settings with 256 nodes. The decline in speed in Fig. 4(b) is attributed to the increased overhead introduced by the compute-bound. While a more complex tree can improve acceleration, it does so at the cost of speed due to intensive matrix multiplications for linear layers and self-attention. The acceleration rate increase follows a logarithmic trend and slows down when the tree size grows as shown in Fig. 4(a). However, the initial gains are substantial, allowing Medusa to achieve significant speedups. If the acceleration increase is less than the overhead, it will slow down overall performance. For detailed study, please refer to Appendix G.
Figure 5: Performance comparison of Medusa using proposed typical sampling. The model is fully fine-tuned from Vicuna-7B. The plot illustrates the acceleration rate and average scores on the writing and roleplay (MT-Bench) with a fixed temperature of 0.7 for 3 different settings: greedy sampling and random sampling (RS) plotted as the star and the dot, and typical sampling curves under different thresholds.
3.3.2 Thresholds of Typical Acceptance
The thresholds of typical acceptance are studied on the writing and roleplay categories from the MT-Bench dataset (Zheng et al., 2023) using Medusa-2 Vicuna 7B. Utilizing the Vicuna 7B model, we aligned our methodology with the approach delineated by  (Hewitt et al., 2022) setting the $\alpha=\sqrt{\epsilon}$. Fig. 5 presents a comparative analysis of our model’s performance across various sampling settings. These settings range from a threshold $\epsilon$ starting at 0.01 and incrementally increasing to 0.25 in steps of 0.01. Our observations indicate a discernible trade-off: as $\epsilon$ increases, there is an elevation in quality at the expense of a reduced acceleration rate. Furthermore, for tasks demanding creativity, it is noted that the default random sampling surpasses greedy sampling in performance, and the proposed typical sampling is comparable with random sampling when $\epsilon$ increases.
@@ 3.3.2 Thresholds of Typical Acceptance
Table 2: Comparison of Different Settings of Vicuna-7B. Quality is obtained by evaluating models on MT-Bench using GPT-4 as the judge (higher the better).
@@ 3.3.3 Effectiveness of Two-stage Fine-tuning
Table 2 shows the performance differences between various fine-tuning strategies for the Vicuna-7B model. Medusa-1, which fine-tunes only the Medusa heads, achieves a 2.18x speedup without compromising generation quality. Medusa-2, which employs two-stage fine-tuning (Section 2.2.2), maintains generation quality and provides greater speedup (2.83x) compared to Medusa-1. In contrast, direct fine-tuning the model with the Medusa heads results in degraded generation quality. The findings indicate that implementing our Medusa-2 for fine-tuning maintains the model’s quality and concurrently improves the speedup versus Medusa-1.
Table 3: Impact of Techniques on Speedup
Technique Speedup
Medusa-1 heads without tree attention $\sim$1.5x
Adding tree attention $\sim$1.9x
Using optimized tree configuration $\sim$2.2x
Training heads with Medusa-2 $\sim$2.8x
4 Discussion
In conclusion, Medusa enhances LLM inference speed by 2.3-2.8 times by equipping models with additional predictive decoding heads, allowing for generating multiple tokens simultaneously and bypassing the sequential decoding limitation. Key advantages of Medusa include its simplicity, parameter efficiency, and ease of integration into existing systems. Medusa avoids the need for specialized draft models. The typical acceptance scheme removes complications from rejection sampling while providing reasonable outputs. Our approach including two efficient training procedures, ensures high-quality output across various models and prompt types. We summarize the development of each technique and their impact on the speedup in Table 3.
In the paper, we focus on the setting with batch size 1 for simplicity. Yet, we want to emphasize that the ideas presented in our paper can be generalized to larger batch-size settings, which are now supported by libraries like TensorRT and Huggingface TGI following our paper.